Find the value of the integral(-p÷2)and (p÷2) v(cos x-cos ³x)dx

Grade:12

1 Answers

Sunil Raikwar
10 years ago
cosx-cos3x=cosx(1-cos2x)=cosx.sin2x{this is an even function}
put sinx=t then cosxdx=dt
so required integral is
2integral{limit 0 to 1}t2dt=2[t3/3]01=2/3

ThanKs & Regards
Sunil Raikwar
askIITians faculty